# Toward Relative Positional Encoding in Spiking Transformers

### v1

@@ 5.4 Patch-based Image Classification
We also conduct experiments on ViT-based SNNs, including Spikformer and the current SOTA model, QKFormer. As discussed in Section 4.5, we extend Gray-PE into a two-dimensional form and integrate it into the patch-based framework to enhance the spatial relationship understanding of spiking Transformers. The results of these experiments are presented in Table 3, where we draw conclusions that:
(1) Gray-PE enhances the performance of Spikformer with less parameters. Gray-PE, whether in 1D or 2D form, consistently improves Spikformer’s classification accuracy. As an approximation of RPE, Gray-PE successfully outperforms spiking absolute PE (i.e., CPG-PE), highlighting the strong dependency between patches within an image. Notably, Spikformer-XNOR with Gray-PE 2D achieves the highest performance across all three datasets, demonstrating the effectiveness of 2D form Gray-PE in capturing relative positional information.
(2) On image classification, QKFormer is insensitive to positional encoding. Surprisingly, QKFormer shows limited improvements or even performance drops when using PE techniques. Whether with CPG-PE or Gray-PE, QKFormer fails to achieve significant performance gains and, in fact, shows a decline in performance. We attribute this phenomenon to the special attention mechanism in QKFormer, known as “q-k attention”, which sums the query matrix along the channel dimension before performing a dot product with key matrix. This strategy eliminates the positional representation provided by both CPG-PE and Gray-PE, which negatively impacts performance.
(3) For Gray-PE, the 2D form performs better than the 1D form on processing image patches. Through a comparison of the performance of SNNs using Gray-PE 1D and 2D, we find that the design of the two-dimensional form of Gray Code is effective. In the case of Spikformers, Gray-PE 2D significantly outperforms Gray-PE 1D, with a $0.44\%$ accuracy improvement averagely. Even for QKFormer, where PE’s influence is limited, Gray-PE 2D still far outperforms its 1D variant. These results confirm that Gray-PE 2D is superior to Gray-PE 1D in handling positional information for image classification tasks, validating the success of our proposed Gray-PE 2D.
5.5 Positional Encoding Analysis
In this section, we will analyze the following points: (1) Impact of the number of bits on the performance of SNNs when using Gray-PE. (2) Rationale behind why we use an approximate RPE rather than a complete one in SNNs.
@@ 5.5 Positional Encoding Analysis
For the first question, we present the results of our parameter sensitivity experiments in Figure 3. Consider that: if the number of bits used for encoding relative positions in Gray-PE is $b$, then the total number of unique encodings possible is $2^{b}$. We set the maximum sequence length is $L$, so relative distances range from $1$ to $L-1$. According to the pigeonhole principle, if $L-1>2^{b}$, there will inevitably be at least two distances that are represented identically. This issue can be mitigated by increasing $b$ to cover the range of relative distances up to $L-1$. From Figure 3 (a), we observe that for long-sequence datasets, such as Solar and Electricity (Length $=168$), the number of bits should be at least $7$ to avoid Gray-PE missing relative positional information. However, for shorter datasets like Metr-la (Length $=12$) and ChnSenti (Length $=32$), $5$ bits are sufficient.
Table 4: Time-series forecasting performance of SNNs trained with Complete RPE (C-RPE) method. We take Metr-la and Electricity as examples. denotes failure to converge.
@@
Regarding the second question, we show the performance of SNNs using complete RPE (C-RPE). As discussed in Section 4.4, C-RPE is obtained by setting $R_{i,j}$ as $\frac{L-1}{|i-j|}$ and adding it to the attention map. However, the large values of this method can cause the vanilla attention map to fail in recording accurate similarity scores. As shown in Table 4, we observe that SNNs with C-RPE significantly underperform compared to other variants, with some even failing to converge, due to the detrimental impact of the large positional encoding values on the training process.
Besides, we design another extended form of Gray-PE, leveraging hierarchical relative positional representation. However, we find that when an encoding method involves carry-over, ensuring the uniqueness of the spike representation for each relative distance becomes challenging. A detailed analysis of this issue can be found in Appendix C.
6 Conclusion
In this paper, we propose a series of relative positional encoding methods for spiking Transformers. Our approach preserves the spiking nature of SNNs while effectively representing relative distances. Experimental evaluations on time series forecasting, text classification, and image classification demonstrate significant performance improvements. These results validate the effectiveness of the proposed RPE method and highlight its potential to enhance the versatility and applicability of SNNs across various domains. Future work and limitations are discussed in Appendix D.
Impact Statement
This research enhances spiking neural networks by introducing an approximation method for relative positional encoding method based on Gray Code. The proposed approach broadens the applicability of SNNs across diverse domains, enabling more robust and versatile neural architectures. Lastly, we do not think our work will negatively impact ethical aspects or future societal consequences.
Appendix A Proof of Theorem 4.1
This section provides a detailed mathematical proof of Theorem 4.1.
A.1 Statement
For any non-negative integer $n$, and for any pair of integers $a$ and $b=a+2^{n}$, the Hamming distance between their Gray Code representations $G(a)$ and $G(b)$ is:
A.2 Proof
We aim to determine the Hamming distance $d_{H}(G(a),G(a+2^{n}))$.
A.2.1 Case 1: $n=0$
When $n=0$, $b=a+1$. By the definition of the Gray Code:
$G(a)=a\oplus\left(a\gg 1\right),\quad G(a+1)=(a+1)\oplus\left((a+1)\gg 1\right).$ (18)
The Gray Code inherently ensures that consecutive integers differ by exactly one bit. Therefore:
$d_{H}(G(a),G(a+1))=1.$ (19)
This establishes the base case for $n=0$.
A.2.2 Case 2: $n\geq 1$
For $n\geq 1$, consider $b=a+2^{n}$. We analyze the XOR of their Gray Codes:
$\begin{split}&G(a)\oplus G(a+2^{n})\\ =&\left(a\oplus\left(a\gg 1\right)\right)\oplus\left((a+2^{n})\oplus\left((a+2^{n})\gg 1\right)\right).\end{split}$ (20)
Rearranging the terms:
$\begin{split}&G(a)\oplus G(a+2^{n})\\ =&\left(a\oplus(a+2^{n})\right)\oplus\left((a\gg 1)\oplus\left((a+2^{n})\gg 1\right)\right).\end{split}$ (21)
Analyzing $a\oplus(a+2^{n})$: Adding $2^{n}$ to $a$ flips the $n$-th bit of $a$. Therefore:
$a\oplus(a+2^{n})=2^{n}.$ (22)
Analyzing $(a\gg 1)\oplus((a+2^{n})\gg 1)$: Shifting both $a$ and $a+2^{n}$ to the right by one bit:
$(a+2^{n})\gg 1=\left(a\gg 1\right)+2^{n-1}.$ (23)
Thus:
$\begin{split}&(a\gg 1)\oplus\left((a+2^{n})\gg 1\right)\\ =&\left(a\gg 1\right)\oplus\left(\left(a\gg 1\right)+2^{n-1}\right)\\ =&2^{n-1}.\end{split}$ (24)
Combining the Results:
$G(a)\oplus G(a+2^{n})=2^{n}\oplus 2^{n-1}.$ (25)
In binary representation, $2^{n}$ and $2^{n-1}$ have single 1’s in distinct positions. Their XOR results in a binary number with exactly two 1’s:
$2^{n}\oplus 2^{n-1}=2^{n}+2^{n-1}.$ (26)
Therefore, the Hamming weight (number of 1’s) of $G(a)\oplus G(a+2^{n})$ is $2$.
Therefore, for $n\geq 1$, since the XOR of the Gray Codes yields a binary number with exactly two 1’s, the Hamming distance between $G(a)$ and $G(a+2^{n})$ is:
$d_{H}(G(a),G(a+2^{n}))=2.$ (27)
A.3 Final Conclusion
Combining both cases, we conclude that for any non-negative integer $n$:
Appendix B Experimental Settings
B.1 Datasets
B.1.1 Time-series Forecasting
@@
We strictly follow the dataset settings of CPG-PE (Lv et al., 2024a). The datasets we used are as follows: Metr-la (Li et al., 2017b): Average traffic speed data collected from highways in Los Angeles County. Pems-bay (Li et al., 2017b): Average traffic speed data from the Bay Area. Electricity (Lai et al., 2018): Hourly electricity consumption data in kilowatt-hours (kWh) of $321$ clients. Solar (Lai et al., 2018): Solar power production. The detailed statistical characteristics and distribution ratios for each dataset are presented below:
Table 5: The statistics of time-series datasets.
@@
B.1.2 Text Classification
For text classification, we follow Lv et al. (2023b) to conduct experiments on six easy discrimination tasks, covering both English and Chinese datasets. Here are the datasets we used in text classification experiments:
The MR dataset, which stands for Movie Review, contains movie-review documents labeled based on their overall sentiment polarity (positive or negative) or subjective rating (Pang & Lee, 2005).
SST-$5$ includes $11,855$ sentences from movie reviews for sentiment classification across five categories: very negative, negative, neutral, positive, and very positive (Socher et al., 2013).
SST-$2$ is the binary version of SST-$5$, containing only two classes: positive and negative. The Subj dataset is designed to classify sentences as either subjective or objective [^1] https://www.cs.cornell.edu/people/pabo/movie-review-data/.
ChnSenti consists of approximately $7,000$ Chinese hotel reviews, each annotated with a positive or negative label [^2] https://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/ChnSentiCorp_htl_all/ChnSentiCorp_htl_all.csv.
Waimai contains around $12,000$ Chinese user reviews from a food delivery platform, intended for binary sentiment classification (positive and negative) [^3] https://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/waimai_10k/waimai_10k.csv.
B.1.3 Image Classification
@@
B.2 Time-series Forecasting
@@
we set the training batch size as $32$ and adopt Adam (Kingma & Ba, 2014) optimizer with a cosine scheduler of learning rate $1\times 10^{-4}$. An early stopping strategy with a tolerance of $30$ epochs is adopted. For other configurations, we honestly follow the SeqSNN framework  [^4] https://github.com/microsoft/SeqSNN proposed by Lv et al. (2024b). We conducted time-series forecasting experiments on 24G-V100 GPUs. On average, a single experiment takes about $1$ hour under the settings above.
B.3 Text Classification

### v2

@@ 5.4 Patch-based Image Classification
Table 3: Accuracy (%) on image classification benchmarks. Numbers with denote our implementations. The best and second-best results are highlighted in bold and underlined formats, respectively. The results with shading are ours. Results are averaged across $4$ random seeds.
Model PE CIFAR10 CIFAR10-DVS CIFAR100 Tiny-ImageNet Avg.
Spike Type Param (M) Acc Param (M) Acc Param (M) Acc Param (M) Acc
Vision-Transformer A $9.32$ $\bm{96.73}$ $9.36$ $\bm{81.02}$ $9.40$ $\bm{62.18}$
Spikformer w/ Conv-PE (Original) A $9.32$   $\;\;94.80^{*}$ $2.57$   $\;\;78.10^{*}$ $9.36$   $\;\;77.04^{*}$ $9.40$   $\;\;48.10^{*}$ $74.51$
Spikformer w/ CPG-PE A $8.17$ $95.06$ $2.06$ $\underline{78.40}$ $8.20$ $77.82$ $8.24$ $\;\;\underline{48.52}^{*}$ $\underline{74.95}$
Spikformer-XNOR w/ Gray-PE 1D R $8.00$ $\underline{95.46}$ $1.99$ $77.90$ $8.04$ $\underline{78.12}$ $8.08$ $48.33$ $\underline{74.95}$
Spikformer-XNOR w/ Gray-PE 2D R $8.00$ $\bf 95.66$ $1.99$ $\bf 78.70$ $8.04$ $\bf 78.45$ $8.08$ $\bf 48.74$ $\bf 75.39$
In this section, we evaluate ViT-based SNNs, Spikformer, which adopts a patch-splitting processing approach. To enhance compatibility with this framework, we extend Gray-PE into a 2D form and integrate it into the patch-based architecture. The experimental results are summarized in Table 3. We draw conclusions that:
(1) Gray-PE enhances the performance of Spikformer while maintaining parameter efficiency. Both 1D and 2D variants of Gray-PE consistently improve classification accuracy. Notably, Gray-PE surpasses spiking absolute PE (CPG-PE), indicating its superior ability to model inter-patch dependencies within images, even as an approximation of RPE.
(2) The 2D variant of Gray-PE demonstrates superior performance over its 1D counterpart in processing image patches. Empirical comparisons between Spikformers equipped with Gray-PE 1D and 2D reveal that the two-dimensional form is highly effective. Specifically, Gray-PE 2D achieves an average accuracy improvement of $0.44\%$ over Gray-PE 1D.
Furthermore, we present the image classification performance of the state-of-the-art QKFormer integrated with our proposed RPE methods in Appendix C.
5.5 Capability of Processing Long Sequences
In this section, we assess the effectiveness of our proposed relative positional encoding methods in handling long sequences within spiking Transformers. To this end, we use two text classification datasets characterized by long input samples: AGNEWS [31] and IMDB [32]. Following [33], we fix the sequence max length to $1024$ for AGNEWS and $2048$ for IMDB. We train the Spikformer model using various positional encoding strategies on these datasets, and present the results in Table 4.
Table 4: Accuracy (%) on $2$ long text classification benchmarks. We set the sentence length to $1024$ for AGNEWS and $2048$ for IMDB.
Model PE AGNEWS IMDB Avg.
Spike Type
Fine-tuned BERT A $\bf 94.50$ $\bf 92.10$ $\bf 93.30$
Spikformer w/ Conv-PE (Original) A $83.84$ $79.08$ $81.46$
Spikformer w/ CPG-PE A $84.70$ $79.47$ $82.09$
Spikformer-XNOR w/ Gray-PE R $84.92$ $79.79$ $82.36$
Spikformer-XNOR w/ Log-PE R $\bf 86.77$ $\bf 80.46$ $\bf 83.62$
As shown in Table 4, although Spikformer models lag behind the fine-tuned BERT in overall performance, both Log-PE and Gray-PE demonstrate effectiveness when handling long input sequences. Notably, Log-PE yields substantial performance improvements, suggesting its strong suitability for processing long texts. This outcome is expected, as Log-PE is specifically designed to accommodate long-range dependencies.
5.6 Analysis and Ablation
In this section, we analyze the following aspects: (1) The influence of internal properties in Gray-PE, (2) Ablation studies on XNOR and Log-PE (shown in Appendix B).
@@ 5.6 Analysis and Ablation
Consider that: If the number of bits used for encoding relative positions in Gray Code is $b$, then the total number of unique encodings possible is $2^{b}$. We set the maximum sequence length is $L$, so relative distances range from 00 to $L-1$. According to the pigeonhole principle, if $L-1>2^{b}$, there will be at least two distances that are represented identically. This issue can be mitigated by increasing $b$ to cover the range of relative distances up to $L-1$. From Figure 3 (a), we observe that for long-sequence datasets, such as Solar and Electricity (Length $=168$), the number of bits should be at least $7$ to avoid Gray-PE missing relative positional information. However, for shorter datasets like Metr-la (Length $=12$) and ChnSenti (Length $=32$), $5$ bits are sufficient.
6 Conclusion
In this work, we have designed several RPE methods for spike Transformers. Our approach preserves the spiking nature of SNNs while effectively representing relative positions. Experimental evaluations on time series forecasting, text classification, and image classification demonstrate significant performance improvements. These empirical results, together with theoretical analysis of the proposed RPE methods, highlight the potential to enhance the versatility and applicability of SNNs across various domains. Future work and limitations are discussed in Appendix E.
Appendix A Proof of Theorem 1
This section provides a detailed mathematical proof of Theorem 1. We use the standard Reflected Binary Code (RBC) $G(x)$ as our Gray Code:
Definition 1.
The Reflected Binary Code (Gray Code) of an integer $x$ is defined as:
$G(x)=x\oplus(x\gg 1),$ (13)
where $\oplus$ denotes the bitwise XOR operation, and $\gg$ denotes the arithmetic right shift.
and we restate Theorem 1 here:
Theorem 1.
For any non-negative integer $n$, and for any pair of decimal integers $a$ and $b=a+2^{n}$, the Hamming distance between their Gray Code representations $G(a)$ and $G(b)$ is consistently:
Here, Hamming distance is the number of different bits between two binary representations.
Proof.
For $n\geq 1$, consider $b=a+2^{n}$. We analyze the XOR of their Gray Codes:
$G(a)\oplus G(a+2^{n})=\left[a\oplus\left(a\gg 1\right)\right]\oplus\left[(a+2^{n})\oplus\left((a+2^{n})\gg 1\right)\right].$ (15)
Using the associativity and commutativity of XOR, we regroup:
$G(a)\oplus G(a+2^{n})=\left[a\oplus\left(a+2^{n}\right)\right]\oplus\left[(a\gg 1)\oplus\left((a+2^{n})\gg 1\right)\right].$ (16)
Let us denote:
$\Delta_{1}=a\oplus\left(a+2^{n}\right),\quad\Delta_{2}=(a\gg 1)\oplus\left(\left(a+2^{n}\right)\gg 1\right).$ (17)
Case 1: $n=0$
In this case, $b=a+1$, and it is well known that adjacent integers in the Gray code differ by exactly one bit. Therefore, we have $d_{H}(G(a),G(b))=d_{H}(G(a),G(a+1))=1$.
Case 2: $n\geq 1$
We consider two subcases based on the bit at position $n$ in $a$.
Subcase A: Bit $n$ in $a$ is 0
Then $a+2^{n}$ flips bit $n$ from 0 to 1, with no carry. Hence:
$\Delta_{1}=2^{n},\quad\Delta_{2}=2^{n-1}.$ (18)
Therefore,
$G(a)\oplus G(b)=2^{n}\oplus 2^{n-1}.$ (19)
This value has exactly two bits set (at positions $n$ and $n-1$), so the Hamming distance is 2.
Subcase B: Bit $n$ in $a$ is 1
Then adding $2^{n}$ to $a$ causes a carry from bit $n$ upwards. Let $c$ be the smallest index greater than $n$ such that bit $c$ in $a$ is 00; bits $n$ through $c-1$ are all $1$. Then:
$\Delta_{1}=a\oplus(a+2^{n})=\sum_{i=n}^{c}2^{i}=\underbrace{1\dots 1}_{c-n+1}\underbrace{0\dots 0}_{n}\ {}_{(2)},$ (20)
has ones at bits $n$ through $c$. We denote $\cdot_{(2)}$ for binary representation. Similarly,
$\Delta_{2}=(a\gg 1)\oplus((a+2^{n})\gg 1)=\sum_{i=n-1}^{c-1}2^{i}=\underbrace{1\dots 1}_{c-n+1}\underbrace{0\dots 0}_{n-1}\ {}_{(2)},$ (21)
has ones at bits $n-1$ through $c-1$. Thus,
$G(a)\oplus G(b)=\Delta_{1}\oplus\Delta_{2}=(\sum_{i=n}^{c}2^{i})\oplus(\sum_{i=n-1}^{c-1}2^{i})=2^{c}+2^{n}$ (22)
has ones at only positions $c$ and $n-1$, and all other bits are canceled due to alignment. The result has exactly two bits set, so the Hamming distance is $2$.
Combining all cases, we conclude that for any non-negative integer $n$:
Appendix B Ablation Study on Log-PE and XNOR
In this section, we compare the performance of vanilla spiking Transformers with Log-PE, XNOR variants with Log-PE, and models equipped with complete relative positional encoding (C-RPE). As discussed in Section 4.4, C-RPE is implemented by setting $R_{i,j}=\frac{L-1}{|i-j|+1}$ and adding it directly to the attention scores.
Table S1: Ablation study on XNOR and Log-PE. We take the time-series forecasting performance of SNNs on Metr-la and Electricity as examples. C-RPE denotes Complete RPE. $\uparrow$ ($\downarrow$) indicates that the higher (lower) the better. denotes failure to converge.
@@ Appendix B Ablation Study on Log-PE and XNOR
As shown in Table S1, both Spikformer and QKFormer with C-RPE perform significantly worse than their Log-PE counterparts, with some variants failing to converge entirely. This degradation is attributed to the overly large positional encodings disrupting the training dynamics. Furthermore, observed from vanilla Spikformer with Log-PE, we confirm that the dot product, which does not use Hamming distance to measure relative distance, is not suitable for RPE methods.
Appendix C Performance of QKFormers on Image Classification
In this section, we conduct experiments on current SOTA spiking Transformer, QKFormer [20].
Table S2: Accuracy (%) of QKFormer on image classification benchmarks. Numbers with denote our implementations. “PE” stands for positional encoding. “R” denotes relative PE, while “A” denotes absolute PE. Results are averaged across $4$ random seeds.
Model PE CIFAR10 CIFAR10-DVS CIFAR100 Tiny-ImageNet Avg.
Spike Type Param (M) Acc Param (M) Acc Param (M) Acc Param (M) Acc
Vision-Transformer A $9.32$ $\bm{96.73}$ $9.36$ $\bm{81.02}$ $9.40$ $\bm{62.18}$
QKFormer w/ Conv PE (Original) $6.74$ $\;\;\underline{96.32}^{*}$ $1.50$   $\;\;\bf 83.40^{*}$ $6.76$   $\;\;\bf 80.90^{*}$ $6.78$   $\;\;\bf 58.07^{*}$ $\bf 79.67$
QKFormer w/ CPG-PE A $7.01$ $96.30$ $1.58$ $82.00$ $7.04$ $80.52$ $7.08$   $\;\;56.75^{*}$ $78.89$
QKFormer-XNOR w/ Gray-PE 1D R $6.02$ $96.22$ $1.41$ $82.20$ $6.04$ $80.48$ $6.06$ $57.21$ $79.03$
QKFormer-XNOR w/ Gray-PE 2D R $6.02$ $\bf 96.36$ $1.41$ $\underline{83.10}$ $6.04$ $\underline{80.82}$ $6.06$ $\underline{57.94}$ $\underline{79.56}$
As shown in Table S2, we find that: QKFormer exhibits insensitivity to positional encoding in image classification. QKFormer exhibits minimal performance gains, or even degradation, when augmented with PE techniques, including both CPG-PE and Gray-PE. We attribute this to its attention design, which aggregates queries along the channel dimension before the dot product with keys. This design inherently biases the model toward spatially specific features while suppressing temporal dependencies. Previous studies [34, 35, 36, 37] have shown that patch-based image classification primarily focuses on spatial (i.e., channel-wise) information rather than the sequential dependencies between patches. This contrasts with sequence modeling tasks such as time-series forecasting and text classification, where capturing inter-token dependencies is crucial. In image classification, our positional encoding encourages the model to emphasize sequential relationships between patches, which introduces a conflict with the QKFormer’s attention mechanism, ultimately hindering performance in this domain.
Appendix D Experimental Settings
D.1 Datasets
D.1.1 Time-series Forecasting
@@
We strictly follow the dataset settings of CPG-PE [7]. The datasets we used are as follows: Metr-la [27]: Average traffic speed data collected from highways in Los Angeles County. Pems-bay [27]: Average traffic speed data from the Bay Area. Electricity [28]: Hourly electricity consumption data in kilowatt-hours (kWh) of $321$ clients. Solar [28]: Solar power production. The detailed statistical characteristics and distribution ratios for each dataset are presented below:
Table S3: The statistics of time-series datasets.
@@
D.1.2 Text Classification
For text classification, we follow [10] to conduct experiments on six easy discrimination tasks, covering both English and Chinese datasets. Here are the datasets we used in text classification experiments:
AGNEWS [31] is a large-scale text classification benchmark derived from AG’s corpus of news articles, containing $120,000$ training samples and $7,600$ valid samples evenly distributed across four categories—World, Sports, Business, and Science/Technology. IMDB [32] is a benchmark for binary sentiment classification, containing $50,000$ movie reviews labeled as positive or negative, split evenly into training and test sets to evaluate natural language understanding and opinion mining models. The MR dataset, which stands for Movie Review, contains movie-review documents labeled based on their overall sentiment polarity (positive or negative) or subjective rating [29]. SST-$5$ includes $11,855$ sentences from movie reviews for sentiment classification across five categories: very negative, negative, neutral, positive, and very positive [30]. SST-$2$ is the binary version of SST-$5$, containing only two classes: positive and negative. The Subj dataset is designed to classify sentences as either subjective or objective***https://www.cs.cornell.edu/people/pabo/movie-review-data/. ChnSenti consists of approximately $7,000$ Chinese hotel reviews, each annotated with a positive or negative labelhttps://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/ChnSentiCorp_htl_all/ChnSentiCorp_htl_all.csv. Waimai contains around $12,000$ Chinese user reviews from a food delivery platform, intended for binary sentiment classification (positive and negative)https://raw.githubusercontent.com/SophonPlus/ChineseNlpCorpus/master/datasets/waimai_10k/waimai_10k.csv.
D.1.3 Image Classification
@@
The Tiny-ImageNet dataset is a simplified subset of the original ImageNet, designed for efficient experimentation in image classification and deep learning research. It consists of $200$ object classes, each containing $500$ training images, $50$ validation images, and $50$ test images (totaling $100,000$ training, $10,000$ validation, and $10,000$ test images). All images are downsampled to a resolution of $64\times 64$ pixels, balancing computational feasibility with visual complexity Designed for efficient deep learning research, it reduces computational costs while maintaining diversity.
@@
D.2 Time-series Forecasting
@@
we set the training batch size as $32$ and adopt Adam [38] optimizer with a cosine scheduler of learning rate $1\times 10^{-4}$. An early stopping strategy with a tolerance of $30$ epochs is adopted. For other configurations, we honestly follow the SeqSNN framework §§§https://github.com/microsoft/SeqSNN proposed by [6]. We conducted time-series forecasting experiments on 24G-V100 GPUs. On average, a single experiment takes about $1$ hour under the settings above.
D.3 Text Classification